# HoneyGPT: Breaking the Trilemma in Terminal Honeypots with Large Language Model

### v1

@@ 5.2.1. Evaluation Metrics
  1. (1)
    Whether the execution of commands succeeds or not. Attack behaviors typically exhibit temporal dependencies and control logic, where the execution of commands relies on the desired effects of previous attack commands, such as successful execution or system information that is of interest to attackers. Thus, a higher execution success rate indicates the honeypot’s capability to induce attackers into deeper interactions and capture more valuable information.
  2. (2)
    Whether the outputs of honeypot comply with the OS logic or not. Whether the honeypot’s outputs align with the logic of the operating system directly affects the effectiveness of its deception. Deviations from the expected behaviors of an operating system, such as inconsistent shell prompts or command execution results, may raise suspicion by attackers and expose the honeypot. A higher level of OS logic compliance indicates a stronger deceptive capability of the honeypot.
@@
Figure 4. Distribution of Deception Categories Across Different Honeypots
Figure 5. Performance Comparison
@@
5.2.2. Performance Comparison
Figures 4 and 5 illustrate the performance comparison among Cowrie, a real operating system, and HoneyGPT based on GPT-3.5 and GPT-4.
In terms of attack success rate and temptation, HoneyGPT based on GPT-3.5 and GPT-4 outperforms both Cowrie and the real operating system. The inherent characteristics of LLMs enable HoneyGPT to better cater to attackers’ intentions, enticing them to further engage in attacks, which is a significant advantage of HoneyGPT. Furthermore, increasing the model parameter size leads to improvements in both metrics. Regarding OS logic compliance and accuracy, HoneyGPT based on GPT-3.5 and GPT-4 also demonstrates superior performance than Cowrie and approaches the level of the real operating system. Additionally, increasing the model parameter size leads to enhancements in both metrics. Note that with the increase in model parameter size, there is a notable improvement in the proportion of successful execution of deceptive commands that comply with the operating system logic (SALC). This indicates that as the model evolves, the precision of HoneyGPT in catering to attackers’ intentions also increases. In summary, the performance comparison highlights the advantages of HoneyGPT based on GPT-3.5 and GPT-4 over Cowrie and the real operating system. HoneyGPT demonstrates superior performance in terms of accuracy, temptation, attack success rate, and compliance with operating system logic. Moreover, increasing the model parameter size can further improve the performance in these metrics.
5.3. Baseline Comparison of Interaction
@@
5.3.1. Performance Comparison
Table 4. Evaluation of Interaction Level
@@ 5.3.1. Performance Comparison
Table 4 presents a detailed comparison of the interaction levels for each honeypot. The results clearly demonstrate the superior performance of HoneyGPT, especially the version based on the GPT-4 API, achieving nearly perfect interaction performance across all categories. This indicates the superiority of HoneyGPT’s capability to maintain interactive sessions and support command interactions.
5.3.2. Unresponsiveness Analysis and Error Origins
@@
Our analysis employs the ”Techniques” component of the ATT&CK framework developed by MITRE to classify the test data (Strom et al., 2018). Within the ATT&CK framework, the ”Techniques” section comprehensively outlines the specific methods adversaries employ to achieve their tactical objectives. By examining the response rate and error causes for each technique, we gain deeper insights into the performance capabilities and potential limitations of HoneyGPT under different API supports.
The test results, as detailed in Appendix B, indicate that HoneyGPT demonstrates excellent support in responding to each attack technique, achieving response rates above 99%. Notably, the performance of GPT-4 surpasses that of GPT-3.5 Turbo. This analysis not only highlights the robustness of HoneyGPT’s design in simulating realistic and responsive interaction, but also underscores the improvements in handling complex security scenarios facilitated by advancements from GPT-3.5 to GPT-4.
Figure 6. Analyse of GPT-3.5 Turbo Non-Responses
Figure 7. Analyse of GPT-4 Non-Responses
Additionally, Figures 6 and  7 provide an analysis of the causes and frequencies of non-responses encountered by each version of HoneyGPT against various ATT&CK techniques. A notable observation is substantial reductions in commonly encountered error types with GPT-4, such as ”response wrong format”, ”wrong command”, and ”prompt length exceeds limit”, in comparison to GPT-3.5 Turbo.
Response wrong format indicates issues with the parsing of responses due to discrepancies between the returned format and the expected format.
@@
5.4. Baseline Comparison of Flexibility
The evaluation of honeypot flexibility focuses on three key aspects: scalability across different systems, adaptability in different system configurations, and integration of additional security analysis features. The comparative evaluation is among HoneyGPT, Cowrie, Honeyd, and real-system honeypot. Cowrie operates as a medium-interaction honeypot that provides interactive terminal emulation, and Honeyd serves as a low-interaction honeypot designed for general purposes. Real-system honeypots, on the other hand, are actual operating systems running on virtual machines.
Table 5. Flexibility in Simulating Different Systems
@@ 5.4. Baseline Comparison of Flexibility
Real System High High Low
@@
5.4.1. Flexibility in Simulating Different Systems
The diversity of operating systems requires traditional honeypots to adapt to various system versions and instruction sets. This adaptation frequently demands independent development efforts for each simulated system. However, the advanced generative capabilities of LLMs offer a groundbreaking solution to overcome these challenges. We evaluate the emulation capabilities of these honeypots across different operating systems, including Ubuntu and CentOS. Table 5 lists the expansion costs, simulation levels, and scalability of each honeypot type. Cowrie, as a medium-interaction honeypot, requires extensive modifications to adapt to different system versions, including updates to system information and the file systems, as well as reconfigurations of the instruction set, and thus the expansion costs are considerably high. Honeyd, as a low-interaction honeypot, requires only minimal adjustments in configuration information to accommodate system version changes, offering high scalability. Real-system honeypots, while achieving high simulation fidelity, require the configuration of new physical environments, which significantly increases the associated costs and reduces overall flexibility. HoneyGPT requires only minimal adjustments to system settings in prompt to simulate different systems. The LLMs can automatically generate file system, instruction set, system user, and other details based on the system settings in prompt. Our approach offers high scalability and a superior level of simulation.
Table 6. Evaluation of Configuration Flexibility for Different Configuration
Honeypot Type Network Configuration File System User Configuration Hardware Settings Service Status
@@ 5.4.1. Flexibility in Simulating Different Systems
Real-system honeypot $\times$ $\times$ $\times$
@@
5.4.2. Flexibility in Simulating Different System Configurations
@@
To assess the flexibility of configuration modifications, we conduct a comparative analysis of Cowrie, Honeyd, real-system Honeypot, and HoneyGPT. We categorize configuration changes into five types: network configuration, file system, user configuration, hardware setting, and service status. If configuration updates can be easily achieved by simple command-line or configuration file modifications, we consider it as supporting flexible modification for that configuration. Table 6 shows that HoneyGPT is the only product allowing flexible modifications across all configuration aspects. Honeyd, Cowrie, and real-system honeypots exhibit limited configuration flexibility.
Traditional honeypots often struggle with modification of system configurations, leading to fixed deployments that can be easily identified by attackers. This rigidity undermines their effectiveness as deception tools. However, the advent of LLMs introduces a novel paradigm. HoneyGPT enables direct modifications to system configurations via prompt adjustments, offering unprecedented flexibility and adaptability, thereby enhancing deception capabilities.
5.4.3. Flexibility of Integration with Security Analysis Capability
In the context of practical operations, honeypots primarily act as collectors of attack data, representing just one facet of comprehensive security measures. Beyond this basic role, the analysis of collected attack data is crucial. Traditional honeypot designs, such as Cowrie, Honeyd, require specialized development efforts, thereby incurring significant overheads to integrate these functionalities.
However, the HoneyGPT’s design circumvents this challenge by leveraging the capabilities of ChatGPT to seamlessly integrate additional functionalities through simple prompts. For example, we can employ ChatGPT to analyze attackers’ behavioral logs, enabling functionalities like security policy alerts. As designed in the prompt manager of HoneyGPT, each interaction with LLMs not only inquires about the outcome of the current command but also asks about the potential impact of the attack command on the operating system.
This innovative approach represents a paradigm shift where a honeypot, by leveraging LLMs, seamlessly integrates security analysis capabilities. The utility of the honeypot is greatly enhanced while reducing the cost and complexity associated with the integration of these auxiliary features. Therefore, HoneyGPT presents a cost-effective, scalable, and efficient approach for future honeypots.
5.5. Field Evaluation
@@
The field evaluation focuses on assessing HoneyGPT’s entrapment efficacy in the real world. We have simultaneously deployed HoneyGPT based on GPT-4 and Cowrie inside the same network for four weeks. HoneyGPT is implemented using the best practice deployment strategies, while Cowrie is configured with its default settings. We detail the evaluation procedure and our main findings below.
5.5.1. HoneyGPT Deployment
Our baseline comparisons demonstrate that HoneyGPT significantly enhances deception, interactivity, and flexibility. However, the economic burden associated with the query costs of commercial LLMs presents practical challenges for deployment. Furthermore, these models also impose restrictions on the quantity and frequency of requests from a single user within a given timeframe, potentially limiting their capacity to handle tasks with high concurrent requests. To optimize the cost-effectiveness of HoneyGPT under the constraints of high query costs and stringent request limitations, we adopt a hybrid deployment strategy. This strategy integrates HoneyGPT with the traditional emulated honeypot approach, leveraging the strengths of both to enhance performance while reducing costs.
In the strategic deployment of HoneyGPT, prudent management of command processing is crucial for optimizing efficacy and cost-efficiency. Those commands that are challenging for LLMs to process, as well as those governed by fixed, simple rules, are delegated to emulated honeypots. In other words, the commands that LLMs typically handle inefficiently, such as binary file reading, and those dependent on time mechanisms (e.g., sleep and uptime), as well as commands following simple rules, such as whoami, are rerouted to emulated honeypots.
Furthermore, resolved interactions from HoneyGPT—comprising terminal prompts and command combinations—are stored in a database. If an attacker’s request corresponds to a stored HoneyGPT query, the recorded response from the database is immediately returned. For new, context-sensitive queries, processing is handled by HoneyGPT to ensure responses are coherent and contextually accurate. Due to the repetitiveness of attacker requests in actual intrusion scenarios, this strategy significantly reduces the query load and congestion on commercial LLMs.
@@ 5.5.1. HoneyGPT Deployment
Moreover, to enhance the appeal of the honeypot to specific adversaries, such as those involved in cryptocurrency mining, the simulated environment within HoneyGPT is configured to mimic a high-performance computing server. Prompts are carefully crafted to align with attackers’ intentions, allowing for the successful execution of attack commands. This is intended to modify attacker behaviors and elicit more prolonged and informative interactions, thereby providing deeper insight into their strategies and potentially uncovering more complex attack techniques.
5.5.2. Comparative Analysis of Attacker Engagement
@@
    Figure 8. Fulfillment of Attacker’s Intent
@@
    Figure 9. Command Support Level
@@
  • Content Rigidity. As illustrated in Figure 10, when attackers engage in reconnaissance of the target system, Cowrie typically provides rigid, easily identifiable honeypot responses. Recognizing these as features of a honeypot, attackers often terminate their interaction prematurely. By contrast, HoneyGPT dynamically generates responses tailored to the honeypot setting outlined in the prompt, effectively camouflaging its honeypot nature and maintaining attacker engagement.
Figure 10. Content Rigidity
This comparative analysis highlights HoneyGPT’s superior capabilities in providing dynamic, contextually relevant content that can significantly sway attacker behaviors, support a broader array of command interactions, and effectively disguise its honeypot identity, thereby extending engagement and potentially exposing more advanced attack strategies.
5.5.3. New Attacks Vector Discovered by HoneyGPT
The deployment of HoneyGPT has unveiled new attack vectors that significantly differ from those captured by the traditional Cowrie honeypot. As detailed in Appendix C, HoneyGPT identifies 11 types of attack actions categorized into six different ATT&CK techniques, highlighting the evolving landscape of cyber threats and the critical role of advanced, adaptive honeypot technologies like HoneyGPT in identifying and analyzing emerging attack vectors.
6. Limitations and Future Work
@@
While the integration of LLMs has enhanced the entrapment capabilities of honeypot systems, there remain several limitations manifested in the following aspects:
  • Context Length Restriction. LLMs are inherently limited by the Context length of prompt. When attackers input and execute lengthy command scripts directly at the terminal, LLMs may fail to respond appropriately. Attempts to limit response lengths within the prompts have proven suboptimal. This issue may necessitate the development of proprietary LLMs tailored for extensive token interactions.
  • Request Rate Limit. Although the use of rule-based matching has reduced the amount of unnecessary commercial LLM requests, commercial models’ limitations on request rates remain a significant obstacle to the deployment of honeypot systems. Professional honeypot models, fine-tuned from open-source LLMs, could potentially address this issue.
@@
  • Ineffectiveness against Fixed Attack Sequences. During field testing, we discovered that some attackers employ predetermined attack sequences to target honeypots, regardless of the honeypot’s ability to respond to requests or the quality of the response content. For such attackers, HoneyGPT is unable to prolong the interaction length by catering to their intent.
@@
  • Prompt-based Attacks on ChatGPT. While no prompt-based attacks against ChatGPT have been discovered thus far, this remains a potential vulnerability that needs to be monitored and addressed (Si et al., 2023; Greshake et al., 2023).
@@
Moving forward, the development of LLMs for honeypot applications should focus on mitigating these limitations. Research efforts may include creating more sophisticated and context-aware models that can handle longer token sequences and have a deeper understanding of complex systems. Furthermore, enhancing models with temporal awareness and addressing request rate limits will be crucial for advancing honeypot technology. It will also be paramount to continually monitor for new types of attacks and adapt the LLMs accordingly to maintain robust security measures. Finally, addressing the overfitting problem may involve implementing more nuanced training strategies that emphasize the correct handling of anomalous or incorrect command inputs.
7. Conclusion

### v2

@@ V-B1 Evaluation Metrics
Whether the execution of commands succeeds or not. In real attack scenarios, responses of honeypots significantly influence the trajectory of attacker behavior. Common impediments such as unsupported tools, insufficient privileges, or the absence of targeted information typically result in the termination of the attack engagements. Therefore, a higher command execution success rate is critical, as it demonstrates the honeypot’s ability to maintain engagement and extract valuable information, serving as a key indicator of its deception effectiveness.
Whether the outputs of the honeypot comply with the OS logic or not. Whether the honeypot’s outputs align with the logic of the operating system directly affects the effectiveness of its deception. Deviations from the expected behaviors of an operating system, such as inconsistent shell prompts or command execution results, may raise suspicion by attackers and expose the honeypot. A higher level of OS logic compliance indicates a stronger deceptive capability of the honeypot.
@@
V-B2 Performance Comparison
Figures 4 illustrate the performance comparison among Cowrie, a real operating system, and HoneyGPT based on GPT-3.5 and GPT-4.
In terms of attack success rate and temptation, HoneyGPT based on GPT-3.5 and GPT-4 outperforms both Cowrie and the real operating system. The inherent characteristics of LLMs enable HoneyGPT to better cater to attackers’ intentions, enticing them to further engage in attacks, as a significant advantage of HoneyGPT. Furthermore, increasing the model parameter size leads to improvements in both metrics. Regarding OS logic compliance and accuracy, HoneyGPT based on GPT-3.5 and GPT-4 also demonstrates superior performance than Cowrie and approaches the level of the real operating system. Additionally, increasing the model parameter size leads to enhancements in both metrics. Note that with the increase in model parameter size, there is a notable improvement in the proportion of successful execution of deceptive commands that comply with the operating system logic (SALC). This indicates that as the model evolves, the precision of HoneyGPT in catering to attackers’ intentions also increases. In summary, the performance comparison highlights the advantages of HoneyGPT based on GPT-3.5 and GPT-4 over Cowrie and the real operating system. HoneyGPT demonstrates superior performance in terms of accuracy, temptation, attack success rate, and compliance with operating system logic. Moreover, increasing the model parameter size can further improve the performance in these metrics.
V-C Baseline Comparison of Interaction
@@
V-C1 Performance Comparison
TABLE IV: Evaluation of Interaction Level
@@ V-C1 Performance Comparison
Table IV presents a detailed comparison of the interaction levels for each honeypot. The results clearly demonstrate the superior performance of HoneyGPT, especially the version based on the GPT-4 API, achieving nearly perfect interaction performance across all categories. This indicates the superiority of HoneyGPT’s capability to maintain interactive sessions and support command interactions.
V-C2 Unresponsiveness Analysis and Error Origins
@@
Our analysis employs the ”Techniques” component of the ATT&CK framework developed by MITRE to classify the test data [27]. Within the ATT&CK framework, the ”Techniques” section comprehensively outlines the specific methods adversaries use to achieve their tactical objectives. By examining the response rate and error causes of each technique, we gain deeper insights into performance capabilities and potential limitations of HoneyGPT under different API supports.
Figure 5: Analyse of GPT-3.5 Turbo Non-Responses
Figure 6: Analyse of GPT-4 Non-Responses
Figure 7: Successful Response Rate with GPT-3.5 Turbo and GPT-4 in HoneyGPT
The test results, as detailed in Figure 7, indicate that HoneyGPT demonstrates excellent support in response to each attack technique, achieving response rates above 99%. Notably, the performance of GPT-4 surpasses that of GPT-3.5 Turbo. This analysis not only highlights the robustness of HoneyGPT’s design in simulating realistic and responsive interactions, but also underscores the improvements in handling complex security scenarios facilitated by advancements from GPT-3.5 to GPT-4.
Additionally, Figures 5 and  6 provide an analysis of the causes and frequencies of non-responses encountered by each version of HoneyGPT against various ATT&CK techniques. A notable observation is substantial reductions in commonly encountered error types with GPT-4, such as ”response wrong format”, ”wrong command”, and ”prompt length exceeds limit”, in comparison to GPT-3.5 Turbo.
Response wrong format indicates issues in response parsing due to discrepancies between the returned format and the expected format.
@@
V-D Baseline Comparison of Flexibility
The evaluation of honeypot flexibility focuses on three key aspects: scalability across different systems, adaptability in different system configurations, and integration of additional security analysis features. The comparative evaluation is among HoneyGPT, Cowrie, Honeyd, and real-system. Cowrie operates as a medium-interaction honeypot that provides interactive terminal emulation, and Honeyd serves as a low-interaction honeypot designed for general purposes. Real-system honeypots, on the other hand, are actual operating systems running on virtual machines.
TABLE V: Flexibility in Simulating Different Systems
@@ V-D Baseline Comparison of Flexibility
real-system High High Low
@@
V-D1 Flexibility in operating system and Configuration
TABLE VI: Evaluation of Configuration Flexibility for Different Configuration
Honeypot Network File System User Settings Hardware Service Status
@@ V-D1 Flexibility in operating system and Configuration
real-system $\times$ $\times$ $\times$
@@
System Simulation: The deployment of HoneyGPT only requires the setting of system version information in the prompt. Leveraging its LLM capabilities and prompt engineering, HoneyGPT can automatically handle commands related to file systems, instruction sets, and system user configurations tailored for different operating systems such as Ubuntu and CentOS. As listed in Table V, this approach not only exhibits high scalability but also delivers an exceptional level of simulation. By contrast, traditional honeypots such as Cowrie and Honeyd still struggle to balance expansion costs with simulation levels, resulting in limited scalability.
@@
Configuration Flexibility: We categorize configuration changes into five types: network configuration, file system, user configuration, hardware setting, and service status. If configuration updates can be easily achieved by simple command-line or configuration file modifications, we consider it as supporting flexible modification for that configuration. As detailed in Table VI, where HoneyGPT allows flexible modifications across all these aspects. Traditional honeypots like Cowrie, Honeyd, and real-system honeypots exhibit limited flexibility, often constrained by fixed deployments that can be easily identified and exploited by attackers.
V-D2 Flexibility of Integration with Security Analysis Capability
In practical operations, honeypots primarily act as collectors of attack data, a single aspect of broader security strategies. Beyond this basic role, the analysis of collected attack data is crucial. Traditional honeypots such as Cowrie and Honeyd require specialized development to integrate these functionalities, leading to significant integration overhead. However, as illustrated in 2, HoneyGPT interacts with LLMs in a unique manner— each interaction not only queries the terminal output but also explores the potential impact of the attack on the OS. This is achieved through the strategic use of prompts designed to seamlessly integrate additional functionalities. This innovative approach signifies a paradigm shift, where honeypots leverage LLMs to integrate security analysis capabilities effortlessly.
V-E Field Evaluation
@@
The field evaluation focuses on assessing HoneyGPT’s entrapment efficacy in the real world. We have simultaneously deployed HoneyGPT based on GPT-4 and Cowrie within the same network for three months. HoneyGPT is implemented using the best practice deployment strategies, while Cowrie is configured with its default settings. We detail the evaluation procedure and our main findings below.
V-E1 HoneyGPT Deployment
@@
To enhance the honeypot’s appeal to specific adversaries, such as those involved in cryptocurrency mining, the simulated environment within HoneyGPT is configured to mimic a high-performance computing server. Prompts are carefully crafted to align with attackers’ intentions, allowing for the successful execution of attack commands. This is intended to modify attacker behaviors and elicit more prolonged and informative interactions, thereby providing deeper insight into their strategies and potentially uncovering more complex attack techniques.
V-E2 Comparative Analysis of Attacker Engagement
@@
Figure 9: Command Support Level
@@
Figure 10: Content Rigidity
@@
Content Rigidity. As illustrated in Figure 10, when attackers engage in reconnaissance of the target system, Cowrie typically provides rigid, easily identifiable honeypot responses. Recognizing these as features of a honeypot, attackers often terminate their interaction prematurely. By contrast, HoneyGPT dynamically generates responses tailored to the honeypot setting outlined in the prompt, effectively camouflaging its honeypot nature and maintaining attacker engagement. This comparative analysis highlights HoneyGPT’s superior capabilities in providing dynamic, contextually relevant content that can significantly sway attacker behaviors, support a broader array of command interactions, and effectively disguise its honeypot identity, thereby extending engagement and potentially exposing more advanced attack strategies.
V-E3 New Attack Vectors Discovered by HoneyGPT
HoneyGPT is able to unveil new attack vectors that differ from those captured by the traditional Cowrie honeypot. Employing the MITRE ATT&CK framework  [27] for analysis, as detailed in Table VII,
HoneyGPT identifies six distinct types of attack actions across five ATT&CK technologies. Specifically, techniques T1082, T1016, and T0842 are classified under Discovery tactics, while T1497 and T1027 fall under Defense Evasion tactics. The Discovery tactics are distinguished by (1) utilization of more tools, including lspci, arp, and tcpdump; (2) more precise querying methods, where attackers integrate information search commands with filtering tools; (3) capability to extract a wider variety of information such as device and network information. Meanwhile, the Defense Evasion tactics are characterized by the execution of sleep commands and the use of simple-to-operate Perl scripts, which are designed to circumvent traditional defense mechanisms.
TABLE VII: New Attack Vectors Discovered by HoneyGPT
ATT&CK Technology Action Intention
System Information Discovery (T1082) Using pipeline operators to integrate system querying commands with filtering tools To extract and display targeted system information (Graphics Card, CPU model) for in-depth analysis and subsequent attacks.
System Network Configuration Discovery (T1016) Use command ”arp -a” to display the system’s ARP table, containing mappings between IP addresses and MAC addresses. Display the system’s ARP table, containing mappings between IP addresses and MAC addresses.
System Network Configuration Discovery (T1016) The nano command is used to view the /etc/resolv.conf file. Access the configuration of the DNS resolver.
Network Sniffing (T0842) Use the tcpdump traffic analysis tool to capture the host’s communication traffic. Sniff the traffic to gain information about the target.
Virtualization / Sandbox Evasion (T1497) Execute the sleep command. Preventing the detection of anomalous behavior and identifying whether the target system is a honeypot.
Obfuscated Files or Information (T1027) Execute Perl scripts that have been obfuscated with base64. Execute while avoiding detection and identification.
VI Limitations and Future Work
@@
While the integration of LLMs has enhanced the entrapment capabilities of honeypot systems, there remain several limitations manifested in the following aspects.
@@
Ineffectiveness against Fixed Attack Sequences. Field testing revealed that some attackers employ predetermined attack sequences to target honeypots, unaffected by the honeypot’s ability to respond to requests or the quality of the response content. For such attackers, HoneyGPT is unable to prolong the interaction length by catering to their intent.
@@
Prompt-based Attacks on ChatGPT. While no prompt-based attacks against ChatGPT have been discovered thus far, this remains a potential vulnerability that needs to be monitored and addressed [26, 7].
@@
Context Length Restriction. When attackers run lengthy command scripts at the terminal, LLMs may not respond accurately. Furthermore, when responses from LLMs become overly extended, attempts to limit these within the prompts have been ineffective. Addressing this issue may call for the development of proprietary LLMs optimized for extensive token interactions.
Request Rate Limit. The deployment strategy detailed in the paper has effectively reduced unnecessary commercial LLM requests, ensuring stable honeypot operations under network attack conditions. Considering the possibility that attackers exploit this vulnerability, the request rate limits of commercial LLMs remain a significant obstacle for honeypot deployment. Professional honeypot models, fine-tuned from open-source LLMs, could potentially address this issue.
Future Work. Moving forward, the development of LLMs for honeypot applications should address existing limitations. Research efforts may include creating more sophisticated and context-aware models that can handle longer token sequences and have a deeper understanding of complex systems. Furthermore, enhancing models with temporal awareness and addressing request rate limits will be crucial to advancing honeypot technology. It will also be paramount to keep monitoring for new types of network attacks and adapt the LLMs accordingly to maintain robust security measures. Finally, addressing the overfitting problem may require more nuanced training strategies that emphasize the appropriate handling of anomalous or incorrect command inputs. Implementing these enhancements will advance the development of sophisticated, high-fidelity LLM-based honeynets. These honeynet systems will accurately simulate complex network interactions, significantly enhancing our ability to deceive and capture more advanced attackers, thus deepening our understanding of evolving cybersecurity threats and attack strategies.
Note that the efficacy of smaller and specialized models in honeypot applications is currently hindered by the absence of a high-quality corpus. These models face significant challenges caused by (1) the complex and varied structure of system commands, (2) the dependence of execution outcomes on external information, and (3) the nuanced requirement for knowledge of both offensive and defensive tactics. To date, there are no effective solutions based on small and specialized models for honeypots. In future work, we plan to address these limitations by developing honeypots based on specialized models, utilizing the data gathered from the field evaluations of this study.
VII Conclusion